Your firm sells hand sanitising gel. Sales have been strong. Your firm sold 1,600,000 units at $3.50 each. The cost of goods sold is $0.95 per unit. Depreciation was $700,000 and interest paid was $450,000. The taxation rate is 30%. At the beginning of the year, fixed (non-current) assets were valued at $7,000,000. At the end of the year, they are valued at $7,900,000. At the beginning of the year, net working capital was $600,000. At the end of the year, net working capital was $670,000. What is your firm's free cash flow?

Select one: a. $504,000 b. $1,732,000 c. $2,301,900 d. $1,531,000 e. $1,660,000
Answer #1

B $ 5,600,000 =1600000*3.5 1,520,000 =1600000*0.95 700,000 A 1 Sales 2 Less: 3 Operating expenses 4 Depreciation 5 6 EBIT 7 L

Hence the correct option is d.$1,531,000

*Please rate thumbs up